3(x-9+4x-1)-2(9-x)=-65
1
Expert's answer
2012-10-29T10:56:17-0400
3(x-9+4x-1)-2(9-x)=-65
3x-27+12x-3-18+2x=-65
17x=-65+48=-17
x=-1